OpenStudy (anonymous):

Please help me answer this: Determine whether each of the following points lie in the unit circle. 1. (3/5, 4/5) 2. (0.8, -0.6) 3. (2 2/3, -1/3) 4. (2, -1) 5. (-8/17, -15/17) Thank you!

Directrix (directrix):

cos (theta) = x/r. The radius of the unit circle is 1. That gives cos(theta) = x. Likewise, on the unit circle sin(theta) = y. Also, recall that sin^2 (theta) + cos^2(theta) = 1. For 1. (3/5, 4/5) Does (3/5)^2 + (4/5)^2 = 1 ? If yes, the point is on the unit circle. If no, then the point is not on the unit circle.

Directrix (directrix):

Note: Recall that the maximum and minimum values of the cosine and sine functions are 1 and -1, respectively. @dsrl26

OpenStudy (shamim):

Need to c drawing @directrix

Directrix (directrix):

I don't see that a drawing would help on this problem if a person knows trig properties. In general, I agree that the role of diagrams in mathematics cannot be overstated. @shamim

OpenStudy (anonymous):

Thank you for you for helping me. I really find it hard. @Directrix @shamim

Directrix (directrix):

Did you get an answer for this: (3/5)^2 + (4/5)^2 ? Don't give up just yet. @dsrl26

OpenStudy (anonymous):

I will try to solve it. @Directrix

Directrix (directrix):

Multiply 3/5 times 3/5. Multiply 4/5 times 4/5. Add what you get for each computation together.

OpenStudy (anonymous):

Oh, I got it! Thank you so much! @Directrix

Directrix (directrix):

What did you get? Do the same here: 2. (0.8, -0.6) Does .8 ^2 + (-.06)^2 = 1?

OpenStudy (anonymous):

Yes. Only numbers 3 and 4 do not lie on the unit circle. Thank you so much, Sir!

OpenStudy (anonymous):

My answer for number 1 is 25/25 or 1. @Directrix

Directrix (directrix):

Correct. So, so far the points on #1 and #2 are on the unit circle. Look at #3 3. (2 2/3, -1/3) The x value of 2 and 2/3 cannot be the x coordinate of a point that lies on the unit circle because the radius of the unit circle is 1. When you square 2 2/3 and add it to the square of -1/3, you will exceed 1.So, that point cannot be on the unit circle.

Directrix (directrix):

Now, it's your turn to decide if this point is on the unit circle. 4. (2, -1) @dsrl26